OpenStudy (anonymous):

if y=8x^2+4x-1 is expressed in the form y=a(x-h)^2+k, where a, h, and k are constants, what is the value of k?

OpenStudy (anonymous):

\(k\) is the second coordianate of the vertex. the first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{4}{2\times 8}=-\frac{1}{4}\] therefore \(k\) is what you get when you replace \(x\) by \(-\frac{1}{4}\)